There are nickels, dimes and quarters in a small piggy bank. It totals $3.40. There are 28 coins. If there are three more dimes than nickels, find the number of each type of coin

Found 2 solutions by stanbon, richwmiller:
Answer by stanbon(75887)

There are nickels, dimes and quarters in a small piggy bank. It totals $3.40. There are 28 coins. If there are three more dimes than nickels, find the number of each type of coin
--------
Equatiions:
d = n + 3
n + d + q = 28
5n+10d+25q = 340
---------
Rearrange::
n - d + 0 = -3
n + d + q = 28
n + 2d+5q = 72
--------------------
Note:: Answers are not whoe numbers.
Cheers,
Stan H.

Answer by richwmiller(17219)

d = n + 3
1*n-1*d+0q=-3
1*n + 1*d + 1*q = 28
5*n+10*d+25*q = 340
Gauss Jordan
1,-1,0,-3
1,1,1,28
5,10,25,340
1,0,0,9
0,1,0,12
0,0,1,7

n = 9,d = 12, q = 7
check
45+120+175=340
340=340
using Determinants
n = 315/35 = 9
d = 420/35 = 12
q = 245/35 = 7